want to know more about magnetic flux!


 

Grade:12

1 Answers

Godfrey Classic Prince
11 years ago

Dear jitesh limbachiya,

Magnetic flux is defined as the number of lines of force passing through a surface normally.

The magnetic flux depends on

(i) the strength of the magnetic field.

(ii) the area of the surface.

(iii) the angle between the magnetic field vector and the area vector.

 If the magnetic field is constant, the magnetic flux passing through a surface of vector S is

 \Phi_B = \mathbf{B} \cdot \mathbf{S} = BS \cos \theta,

where B is the magnitude of the magnetic field (the magnetic flux density) having the unit of Wb/m2 (Tesla).

 

Hope this helped you immensely..!

All the Very Best & Good Luck to you ..!!

 

Best Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

Please approve my answer if you liked it by clicking on "Yes" given below..!!Smile